OpenStudy (anonymous):

A spring BB gun fires a BB at a speed of 35 m/s. If the barrel is 1.5 m long, what is the spring constant of the gun?

OpenStudy (anonymous):

Better asked in Physics, but the relevant equations are F=k*s v^2-(v0)^2 = 2a*x and maybe something else. Do you happen to know the mass of the bb and the length of the spring?

OpenStudy (anonymous):

No it was not given, unfortunately.

OpenStudy (anonymous):

Hmm, maybe this can be done in terms of energy . . Energy of a spring is 0.5kx^2 The kinetic energy of the bb is 0.5mv^2 You can probably combine the equations for force and acceleration with those for energy to solve. Are we to assume that the spring moves the entire 1.5m barrel length to fire the bb?

OpenStudy (anonymous):

Yes the entire length

OpenStudy (anonymous):

Ok, then that makes things easier. Would still be better if we knew the mass of the bb. They come in standard sizes, we could assume standard bb gun sized bb to use for the mass... The formula I derived from conservation of energy is k=(v^2/x^2)*m